The current of a river is 5 miles per hour. A boat travels between two points on the river. If it takes 2 hours to travel in one direction and 5/2 hours to travel in the other direction, what is the speed of the boat in still water?
Answer by stanbon(75887)

The current of a river is 5 miles per hour. A boat travels between two points on the river. If it takes 2 hours to travel in one direction and 5/2 hours to travel in the other direction, what is the speed of the boat in still water?
-------
Downstream DATA:
distance = x miles ; time = 2 hrs ; rate = x/2 mph
------
Upstream DATA:
distance = x miles ; time = 5/2 hrs ; rate = x/(5/2) mph
------
Equation:
Down:: b + 5 = x/2
Up:::: b - 5 = (2x)/5
-------
Subtract and solve for "x"::
(x/2)-(2x/5) = 10
Multiply thru by 10 to get:
5x - 4x = 100
x = 100 miles
----
Ans: Solve for "b"::
b + 5 = x/2
b + 5 = 50
boat speed = 45 mph
=======================
Cheers,
Stan H.
-----------------